Circular No. 6419

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

X-RAY TRANSIENT IN THE LARGE MAGELLANIC CLOUD F. E. Marshall, Laboratory for High Energy Astrophysics, Goddard

Space Flight Center (LHEA/GSFC); and R. Corbet and T. Strohmayer, Universities Space Research Association and LHEA/GSFC, report the discovery of a new hard-x-ray source in the direction of the LMC using the RXTE’s Proportional Counter Array. The 2- to 10-keV flux on June 8.19 and 11.31 UT was ∼ 18 mCrab (luminosity 1.3 × 10^{38} erg/s at the distance of the LMC). The best-fit position is α = 5^{h}22^{m}.5 ± 0^{m}.2, δ = −69^{o}37′ ± 9′ (equinox 2000.0). An absorption event on June 11.3154 lasting ∼ 100 s suggests a high-inclination binary system. No pulsations were seen, with an upper limit on the pulsed fraction of 1 percent for frequencies between 0.1 and 40 Hz. Additional x-ray observations are urged to establish an accurate position and possibly to determine a binary period.

ROSAT SOFT-X-RAY MAPS S. L. Snowden, M. J. Freyberg, P. P. Plucinsky, J. H. M. M. Schmitt,

J. Trümper, and W. Voges, Max-Planck-Institut für Extraterrestrische Physik, Garching; R. J. Edgar, D. McCammon, and W. T. Sanders, University of Wisconsin, Madison, report: “The first maps of the soft x-ray diffuse background (SXRB) from the ROSAT XRT/PSPC All-Sky Survey have been released and are available through the World Wide Web (URL

http://www.rosat.mpe-garching.mpg.de/survey/sxrb/) and via anon-
ymous ftp (host ftp.rosat.mpe-garching.mpg.de, directory archive/

survey/sxrb). A detailed description of the data and maps has appeared in Ap.J. 454, 643 (1995). The maps cover ∼ 98 percent of the sky in the bands 0.25, 0.75, and 1.5 keV, with ∼ 2^{o} angular resolution and high sensitivity for low-surface-brightness, extended features. The effects of nonx-ray contamination and x-rays of solar-system origin have been eliminated to the greatest possible extent, but discrete x-ray sources have not been removed. The much-improved angular resolution, statistical precision, and completeness of coverage of these maps reveal considerable structure over the entire energy range 0.1–2.0 keV that was not observed previously. The data compare well with previous all-sky surveys in terms of absolute normalization and zero point. For each energy band, an intensity map, as well as an exposure map, is available in FITS format. All maps were constructed using a Hammer-Aitoff equal-area projection in zero-centered galactic coordinates with longitude increasing to the left. Questions or comments may

be directed to sxrb@rosat.mpe-garching.mpg.de.”

1996 June 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green